# New affine invariant ensemble samplers and their dimensional scaling

### v1

@@ Abstract.
We introduce some new affine invariant ensemble samplers that are easy to construct and improve upon existing widely used algorithms, especially for high-dimensional problems. Specifically, we propose a derivative-free ensemble side move sampler that performs favorably compared to popular samplers in the emcee package. Additionally, we develop a class of derivative-based ensemble Hamiltonian Monte Carlo (HMC) samplers with affine invariance, which outperform standard HMC without affine invariance when sampling highly skewed distributions. We provide asymptotic scaling analysis for high-dimensional Gaussian targets to further elucidate the properties of these affine invariant ensemble samplers. In particular, with derivative information, the affine invariant ensemble HMC can scale much better with dimension compared to derivative-free ensemble samplers.
@@ 1. Introduction
Affine invariant ensemble samplers have proven effective for routine Bayesian inference applications [20], but they may perform poorly in high-dimensional problems [30, 7]. This particularly applies to derivative-free samplers such as the stretch move sampler [26], widely used via the emcee package [20]. Indeed, compelling arguments suggest that “ensemble methods are doomed to fail in high dimensions” [7], based on insightful observations about typical sets: high-dimensional distributions typically concentrate on thin shells, and the interpolation or extrapolation between two points in the stretch move—as well as in many other ensemble samplers—is unlikely to fall within this shell. As a result, small step sizes must be used, and the samplers effectively “devolve into random walks with poorly biased directional choices.”
@@ 1.2. Related work on affine invariant samplers
Among unbiased derivative-based MCMC samplers, HMC achieves the state-of-the-art scaling with dimension [3]. However, HMC often struggles with anisotropic or ill-conditioned target distributions and typically requires adaptation of the mass matrix to correct scale imbalances, as it is not inherently affine invariant. Metrics based on Riemannian manifolds and Hessians have been used to make HMC affine invariant [25, 32, 31], but these approaches generally require costly second-order information. We believe that systematically developing affine invariant ensemble HMC methods — based solely on first-order information — represents a promising algorithmic opportunity to advance the state of ensemble samplers.
@@ 4.3. Hamiltonian walk move sampler
The algorithm may be interpreted as a gradient-enhanced walk move. In one form of the walk move [26], the proposal is $\mathbf{x}_{i}+B_{S^{(1)}}\mathbf{z}$ with $\mathbf{z}\sim\mathcal{N}(0,I_{N/2\times N/2})$. In our algorithm, the derivative of $\mathbf{x}_{i}$ is similarly a linear combination of the columns of $B_{S^{(1)}}$. When the integration time is small, the leading order term follows the same direction as $B_{S^{(1)}}\mathbf{z}$. With longer integration times, our approach adaptively determines the move strength using Hamiltonian-type dynamics that incorporate derivative information of the potential. For this reason, we term the method the Hamiltonian walk move sampler.
@@ 4.4. Hamiltonian side move sampler
In the last subsection, the use of empirical covariance or centered ensembles makes the algorithm affine invariant. This approach accounts for global statistics through covariance. Alternatively, we can use the local side move direction to derive an affine invariant algorithm as follows.
@@
This approach only needs to calculate directional gradients rather than full gradients. We refer to the algorithm as the Hamiltonian side move sampler since it uses precisely the same direction as the standard side move, with the move length adaptively determined by Hamiltonian-type dynamics. Like our previous methods, this approach maintains affine invariance for the same underlying reasons.
@@ 4.5. Analysis of high dimensional scaling
    where $\mathcal{N}(\alpha^{4}\mu_{\rho},\alpha^{4}\sigma_{\rho})$ is a Gaussian distribution with mean and variance
@@ 5.2. Synthetic: Gaussian
We run $2\times 10^{5}$ iterations of these samplers as burn-in, followed by $10^{6}$ steps treated as the stationary phase. Due to memory constraints, we thin the samples by a factor of $10$, compute the autocorrelation function and time for these thinned samples, and then scale the results to approximate the autocorrelation function and time of the original samples. The autocorrelation function is computed for the observable $f(\mathbf{x})=x_{1}$, the first component.
@@
Figure 3. Scaled autocorrelation functions for sampling anisotropic Gaussian targets with condition number $\kappa=1000$. Left: stretch move; right: side move. Scaled lag $=$ original lag///dim×$\times 4$.
For $d=128$, we report the detailed results to one decimal place in Table 1. We observe that the side move achieves approximately $1/2$ of the autocorrelation time of the stretch move. Moreover, with derivative information, samplers become more scalable in high dimensions, as demonstrated by the substantial decrease in autocorrelation time for HMC. We note that when $n=2$, the step size is too large for HMC, causing negligible acceptance rates. However, this large step size works for the affine invariant Hamiltonian walk and side moves. Notably, the Hamiltonian walk move with $n=2$ yields a small autocorrelation time of $12.7$. Compared to HMC, the computational cost is approximately $1/3$, resulting in a speedup of roughly $67.8/12.7\times 3\approx 16$ times. If gradient and function evaluations have comparable costs, then the Hamiltonian walk move with $n=2$ achieves $1000.1/12.7\times 1/4\approx 19.7$ times acceleration compared to the side move.
@@
acceptance autocorrelation func eval grad eval
rate time $\tau_{e}$ or $\tau_{s}$ per iter per iter
Stretch move 0.45 2043.6 1 0
Side move 0.45 1000.1 1 0
HMC: $n=10$ 0.57 67.8 1 11
@@
Hamiltonian walk move: $n=10$ 0.98 10.5 1 11
Hamiltonian walk move: $n=2$ 0.61 12.7 1 3
Hamiltonian side move: $n=10$ 1.00 898.2 1 11
Hamiltonian side move: $n=2$ 0.98 732.3 1 3
Table 1. Performance for anisotropic Gaussian targets with condition number $\kappa=1000$. Dimension $d=128$. For HMC and affine invariant HMC (Hamiltonian walk and side moves), the total integration time is $T=1$ and $n$ is the number of leapfrog iterations.
@@
Figure 4. Autocorrelation time versus dimension for anisotropic Gaussian targets with condition number $\kappa=1000$. For HMC and affine invariant HMC (Hamiltonian walk and side moves), the total integration time is $T=1$ and $n$ is the number of leapfrog iterations.
@@ 5.3. Synthetic: Rings
acceptance autocorrelation func eval grad eval
rate time $\tau_{e}$ or $\tau_{s}$ per iter per iter
Stretch move 0.29 2435.4 1 0
Side move 0.45 355.4 1 0
HMC: $n=10$ 0.69 20.8 1 11
@@
Hamiltonian walk move: $n=10$ 0.99 10.7 1 11
Hamiltonian walk move: $n=2$ 0.72 11.9 1 3
Hamiltonian side move: $n=10$ 1.00 354.8 1 11
Hamiltonian side move: $n=2$ 0.98 309.7 1 3
Table 2. Performance for ring shaped distributions: dimension $d=50$. For HMC and affine invariant HMC (Hamiltonian walk and side moves), the total integration time is $T=1$ and $n$ is the number of leapfrog iterations.
@@ 5.4. Invariant distribution to stochastic PDEs
Figure 5. Autocorrelation time versus dimension for the bimodal SPDE targets. For HMC and affine invariant HMC (Hamiltonian walk and side moves), the total integration time is $T=1$ and $n$ is the number of leapfrog iterations.
acceptance autocorrelation func eval grad eval
rate time $\tau_{e}$ or $\tau_{s}$ per iter per iter
Stretch move 0.44 3021.3 1 0
Side move 0.44 1398.3 1 0
@@
Hamiltonian walk move: $n=10$ 0.98 11.2 1 11
Hamiltonian walk move: $n=2$ 0.59 14.9 1 3
Hamiltonian side move: $n=10$ 1.00 902.8 1 11
Hamiltonian side move: $n=2$ 0.98 770.8 1 3
Table 3. Performance for the bimodal distribution which is the invariant distribution to stochastic PDEs: dimension $d=128$. For HMC and affine invariant HMC (Hamiltonian walk and side moves), the total integration time is $T=1$ and $n$ is the number of leapfrog iterations.
@@ 6. Discussions and Conclusions
Algorithmically, adapting the step size and, in the case of affine invariant HMC, the number of leapfrog steps could further improve performance, according to the wide success of the No-U-Turn Sampler (NUTS) [28, 8]. In our experiments, we observe that the Hamiltonian side move leads to smaller autocorrelation time compared to the standard side move. In each iteration, they move in the same direction, suggesting potential for further adaptation of the move length of the derivative-free side move. A recent relevant development is the No-Underrun Sampler (NURS) [5], which focuses on adapting move lengths along random line directions without using derivatives.

### v2

@@ Abstract.
We introduce new affine invariant ensemble samplers that are easy to construct and improve upon existing algorithms, especially for high-dimensional problems. Specifically, we propose a derivative-free ensemble side move sampler that performs favorably compared to popular samplers in the emcee package. Additionally, we develop a class of derivative-based ensemble Hamiltonian Monte Carlo (HMC) samplers with affine invariance, which outperform standard HMC without affine invariance when sampling highly skewed distributions. We provide asymptotic scaling analysis for high-dimensional Gaussian targets to further elucidate the properties of these affine invariant ensemble samplers. In particular, with derivative information, the affine invariant ensemble HMC can scale much better with dimension compared to derivative-free ensemble samplers.
@@ 1. Introduction
Affine invariant ensemble samplers have proven effective for routine Bayesian inference applications [20], but they may perform poorly in high-dimensional problems [32, 7]. This particularly applies to derivative-free samplers such as the stretch move sampler [26], widely used via the emcee package [20]. Indeed, compelling arguments suggest that “ensemble methods are doomed to fail in high dimensions” [7], based on insightful observations about typical sets: high-dimensional distributions (such as Gaussians) typically concentrate on thin shells, and the interpolation or extrapolation between two points in the stretch move—as well as in many other ensemble samplers—is unlikely to fall within this shell. As a result, small step sizes must be used, and the samplers effectively “devolve into random walks with poorly biased directional choices.”
@@ 1.2. Related work on affine invariant samplers
Among unbiased derivative-based MCMC samplers, HMC [41] achieves the state-of-the-art scaling with dimension [3]. However, HMC can struggle with anisotropic or ill-conditioned target distributions and requires adaptation of the mass matrix [8] to correct scale imbalances, as it is not affine invariant. Metrics based on Riemannian manifolds and Hessians have been used to make HMC affine invariant [25, 34, 33], but these approaches generally require costly second-order information. We believe that systematically developing affine invariant ensemble HMC methods — based solely on first-order information — represents a promising algorithmic opportunity to advance the state of ensemble samplers. We note that researchers have explored using ensemble chains to form a diagonal preconditioner based on individual variance [30], which leads to scale invariance — a property slightly weaker than affine invariance. Related work has also focused on tuning-free multichain adaptive HMC methods that utilize parallel hardware [28, 52, 51].
@@ 4.3. Hamiltonian walk move sampler
The algorithm may be interpreted as a gradient-enhanced walk move. In one form of the walk move [26], the proposal is $\mathbf{x}_{i}+B_{S^{(1)}}\mathbf{z}$ with $\mathbf{z}\sim\mathcal{N}(0,I_{N/2\times N/2})$. In our algorithm, the derivative of $\mathbf{x}_{i}$ is similarly a linear combination of the columns of $B_{S^{(1)}}$. When the integration time is small, the leading order term follows the same direction as $B_{S^{(1)}}\mathbf{z}$. With longer integration times, our approach adaptively determines the move strength using Hamiltonian-type dynamics that incorporate derivative information of the potential. For this reason, we term the method the Hamiltonian walk move sampler. In general, similar to the walk move sampler, we can choose an arbitrary subset $S$ from the complementary ensemble to form the centered ensemble; in the above Hamiltonian walk move algorithm, $S$ is chosen to contain all particles in the complementary ensemble. Choosing a smaller $S$ could reduce the arithmetic cost.
@@ 4.4. Hamiltonian side move sampler
In the last subsection, the use of empirical covariance or centered ensembles of all particles in the complementary ensemble makes the algorithm affine invariant. This approach accounts for global statistics through covariance. Alternatively, we can use the local side move direction to derive an affine invariant algorithm as follows.
@@
This approach only needs to calculate directional gradients rather than full gradients. We refer to the algorithm as the Hamiltonian side move sampler since it uses precisely the same direction as the standard side move, with the move length adaptively determined by Hamiltonian-type dynamics. Like our previous methods, this approach maintains affine invariance for the same underlying reasons. Similar to discussions at the end of Section 3.3, the Hamiltonian side move may be interpreted as a specific local variant (i.e., $|S|=2$) of the Hamiltonian walk move.
@@ 4.5. Analysis of high dimensional scaling
    where $\mathcal{N}(\alpha^{4}\mu_{\rho},\alpha^{4}\sigma_{\rho})$ is a Gaussian distribution with
@@ 5.2. Synthetic: Gaussian
We run $2\times 10^{5}$ iterations of these samplers as burn-in, followed by $10^{6}$ steps treated as the stationary phase. The autocorrelation function [^1] Due to memory constraints, we often thin the samples by a factor of $10$, compute the autocorrelation function and time for these thinned samples. We explicitly note this whenever the thinning is used. is computed for the observable $f(\mathbf{x})=x_{1}$, the first component.
@@
Figure 3. Rescaled autocorrelation functions for sampling anisotropic Gaussian targets with condition number $\kappa=1000$. Left: stretch move; right: side move. Rescaled lag $=$ original lag///dim×$\times 4$.
For $d=128$, we report the detailed results to one decimal place in Table 1. We observe that the side move achieves approximately $1/2$ of the autocorrelation time of the stretch move. Moreover, with derivative information, samplers become more scalable in high dimensions, as demonstrated by the substantial decrease in autocorrelation time for HMC. We note that when $n=2$, the step size is too large for HMC, causing negligible acceptance rates. However, this large step size works for the affine invariant Hamiltonian walk and side moves. Notably, the Hamiltonian walk move with $n=2$ yields a small autocorrelation time of $1.27$. Compared to HMC, the computational cost is approximately $1/3$, resulting in a speedup of roughly $6.78/1.27\times 3\approx 16$ times. If gradient and function evaluations have comparable costs, then the Hamiltonian walk move with $n=2$ achieves $100.01/1.27\times 1/4\approx 19.7$ times acceleration compared to the side move.
@@
acceptance autocorrelation time func eval grad eval
rate (thinning by 10) per iter per iter
Stretch move 0.45 204.36 1 0
Side move 0.45 100.01 1 0
HMC: $n=10$ 0.57 6.78 1 11
@@
Hamiltonian walk move: $n=10$ 0.98 1.05 1 11
Hamiltonian walk move: $n=2$ 0.61 1.27 1 3
Hamiltonian side move: $n=10$ 1.00 89.82 1 11
Hamiltonian side move: $n=2$ 0.98 73.23 1 3
Table 1. Performance for anisotropic Gaussian targets with condition number $\kappa=1000$. Dimension $d=128$. For HMC and affine invariant HMC (Hamiltonian walk and side moves), the total integration time is $T=1$ and $n$ is the number of leapfrog steps.
@@
Figure 4. Autocorrelation time (thinning by $10$) versus dimension for anisotropic Gaussian targets with condition number $\kappa=1000$. For HMC and affine invariant HMC (Hamiltonian walk and side moves), the total integration time is $T=1$ and $n$ is the number of leapfrog steps.
@@ 5.3. Synthetic: Rings
acceptance autocorrelation time func eval grad eval
rate (thinning by 10) per iter per iter
Stretch move 0.29 243.54 1 0
Side move 0.45 35.54 1 0
HMC: $n=10$ 0.69 2.08 1 11
@@
Hamiltonian walk move: $n=10$ 0.99 1.07 1 11
Hamiltonian walk move: $n=2$ 0.72 1.19 1 3
Hamiltonian side move: $n=10$ 1.00 35.48 1 11
Hamiltonian side move: $n=2$ 0.98 30.97 1 3
Table 2. Performance for ring shaped distributions: dimension $d=50$. For HMC and affine invariant HMC (Hamiltonian walk and side moves), the total integration time is $T=1$ and $n$ is the number of leapfrog steps.
@@ 5.4. Invariant distribution to stochastic PDEs
Figure 5. Autocorrelation time (thinning by $10$) versus dimension for the bimodal SPDE targets. For HMC and affine invariant HMC (Hamiltonian walk and side moves), the total integration time is $T=1$ and $n$ is the number of leapfrog steps.
acceptance autocorrelation time func eval grad eval
rate (thinning by $10$) per iter per iter
Stretch move 0.44 302.13 1 0
Side move 0.44 139.83 1 0
@@
Hamiltonian walk move: $n=10$ 0.98 1.12 1 11
Hamiltonian walk move: $n=2$ 0.59 1.49 1 3
Hamiltonian side move: $n=10$ 1.00 90.28 1 11
Hamiltonian side move: $n=2$ 0.98 77.08 1 3
Table 3. Performance for the bimodal distribution which is the invariant distribution to stochastic PDEs: dimension $d=128$. For HMC and affine invariant HMC (Hamiltonian walk and side moves), the total integration time is $T=1$ and $n$ is the number of leapfrog steps.
@@ 6. Discussions and Conclusions
Algorithmically, adapting the step size and, in the case of affine invariant HMC, the number of leapfrog steps could further improve performance, according to the wide success of the No-U-Turn Sampler (NUTS) [29, 8]; in particular, there has been some recent developments on tuning-free multichain adaptive HMC methods [30, 28, 52, 51]. In our experiments, we observe that the Hamiltonian side move leads to smaller autocorrelation time compared to the standard side move. In each iteration, they move in the same direction, suggesting potential for further adaptation of the move length of the derivative-free side move. A recent relevant development is the No-Underrun Sampler (NURS) [5], which focuses on adapting move lengths along random line directions without using derivatives.